4. Find a QR-factorization of the matrix 5. Find an LU-decomposition of the matrix A =
Answer #1

soln 4) To find QR-factorization of the matrix. we find QR-factorization by using Gram-Schmidt process. A=Q R where Q is orth= (2,-2)-(是),) - (2,-2)-(-) Uz = (16) log ( 5) | CS)-(台) 三、主(兰,) = (s) 區(台) = . - ( 台) e = ( 6) | avel = (-1.3)-(-13) aven =:.Q = [eid ez] , R= [ave, ale hal which is or matrix. which upper tringular matrix. A- 57 To find LU Decomposition of given :107 -27 Loin-1- 21 Rat BRI I JA={: 2] - =) ie. W= (112) Now to find L Let E. =3 ] 15= 1 #0 in Ei is invertible. Eq? become. E.: A = Eiu ILU Arlon - L3 I Lo